what velocity in m/s does a falling object acquire after falling for 0.50 s, 1.0 vs, 2.0 s, 5.0 s, and 12 s?
Answer by ikleyn(52858)

.

The general formula is


    u = g*t,


where u is the velocity of the body freely falling (in meters per second), t is the time in seconds, g is the gravity acceleration g = 9.81 m/s^2.


For example, for t = 2 seconds (one of the examples from list)


    u = 9.81*2 = 19.62 m/s.


For other cases, use the formula and make calculations  ON  YOUR  OWN  by the same way,  as instructed.

Solved, answered and explained.